OpenStudy (anonymous):

Rewrite with only sin x and cos x cos 3x

OpenStudy (anonymous):

the choices cos x - 4 cos x sin2x -sin3x + 2 sin x cos x -sin2x + 2 sin x cos x 2 sin2x cos x - 2 sin x cos x

OpenStudy (anonymous):

i thought it was the third one

OpenStudy (anonymous):

please help i don't know if im right

OpenStudy (anonymous):

\(\cos 3x=\cos\left(2x+x\right)\)

OpenStudy (anonymous):

\[\cos(\alpha+\beta)=\cos \alpha \cos \beta-\sin \alpha \sin \beta\]

OpenStudy (anonymous):

i really bad at math i don't understand was my answer choice wrong

OpenStudy (anonymous):

i don't know, i didn't work it out. that's your job

OpenStudy (anonymous):

can you explain how to do it

OpenStudy (anonymous):

use the formula above... then see where it takes you

OpenStudy (anonymous):

i got cos x - 4 cos x sin2x thanks

OpenStudy (anonymous):

wait i think i did it wrong i re-did it i got -sin3x + 2 sin x cos x

OpenStudy (anonymous):

know im just confused

OpenStudy (anonymous):

*now*

OpenStudy (anonymous):

someone please help me

OpenStudy (anonymous):

pgpilot326?

OpenStudy (anonymous):

not gettting your choices... could be using difference formula: cos(3x) = cos(4x-x)

OpenStudy (anonymous):

umm idk what r u getting

OpenStudy (anonymous):

wait... \[\cos \left( 3x \right)=\cos \left( 2x+x \right)=\cos \left( 2x \right)\cos \left( x \right)-\sin \left( 2x \right)\sin \left( x \right)\]\[=\left[ 1-2\sin^2\left( x\right)\right]\cos \left( x\right) -\left[2\sin \left( x \right) \cos\left( x\right)\right]\sin\left( x \right)\]\[=\cos \left( x \right)-4\sin^2\left( x \right)\cos \left( x \right)\]

OpenStudy (anonymous):

its almost like the 2nd answer i got..istill dont get it

OpenStudy (anonymous):

it's the first choice if what you mean by sin2x is \(\sin^2 x\)